Figure 1,2: Left Hand Plot: The optical-infrared (Keck-UKIRT) spectrum of the T dwarf Gliese 229B showing the peak of the SED in the near infrared. Note the extreme red optical-IR colour; I−J=5.8 but blue IR colours; J−H Right Hand Plot: The 0.4-2.5 micron spectrum of the Halo white dwarf WD0346+246 showing the dramatic effects of collision induced absorption by H_{2} in the infrared (Hodgkin et al. 1999). Thus the object appears red in the optical, but blue in the infrared. Two blackbody fits are overlaid at T=3850 K (solid curve) and T=1500 K (dashed curve). We derive T_{e}=3500 K for this object. Photometry is from UKIRT (IR) and the JKT (optical), whilst spectoscopy is from the WHT (optical) and the KECK (infrared).

Figure 3,4; Left Hand Plot: From Storrie-Lombardi, McMahon & Irwin(1996a): The mean cosmological mass density in neutral HI, Ω_{HI}, contributed by damped Lyman-α absorbers for 0.01≤z≤4.7 for q_{0}=0.5. The region Ω_{star} is the ±σ range for the mass density in stars in nearby galaxies (Gnedin & Ostriker 1992). The point at z=0 is the local value (Fall & Pei 1993; Rao & Briggs 1993). Right Hand Plot: Simulated optical and near-IR colour-colour plot showing the locus of quasars compared with galactic stars from Gunn & Stryker 1983. Unlike the conventional purely optical colour-colour plots the redddening vector does not drive the quasars into the stellar locus and hence we have a reddening independent technique for discovering quasars.

References

Fall, S.M. & Pei, Y.C. 1993, ApJ, 402, 479 Fall, S.M. & Pei, Y.C., 1995, ApJ, 454, 69 Gnedin, N.Y. & Ostriker, J.P. 1992, ApJ, 400, 1 Gunn & Stryker 1983 ApJS 52 121 Mobasher et al. 1986 MNRAS 223 11 Burgasser et al., 1999, ApJ, 522, L65 Cuby, J.G., et al., 1999, A&A, in press (astroph/9907028). Hansen, B.M.S., 1998, Nature, 394, 860.

Hodgkin, S.T. etal,, 1999, Nature, submitted (email sth@ast.cam.ac.uk for a preprint). Reid, I.N., et al., 1999, ApJ, in press (astro-ph/9905170) Richer, H.B., Hansen, B.M.S., Limongi, M., Chieffi, A., Straniero, O., Fahlman, G.G., 1999, ApJ, in press (astroph/9908337). Storrie-Lombardi, L.J., McMahon, R.G., Irwin, M. , 1996a, MN-RAS, 283, L79-L83. Storrie-Lombardi, L.J., Irwin, M.,, McMahon, R.G., 1996b, MN-RAS, 282, 1330-1342. Warren, S.J., Hewett, P.C., Foltz, C.B., 2000,MNRAS, 312, 827

PATT2 Version 12/96 Particle Physics and Astronomy Research Council Polaris House, North Star Avenue, Swindon, SN2 1SZ Telephone 0793 442000  Fax 0793 442002 APPLICATION FOR TELESCOPE TIME (OPTICAL AND INFRARED) Refer to Notes for Guidance on the completion of the application form PATT2; type entries in black 1 TELESCOPE Specify one of AAT,WHT,INT,JKT,UKIRT 2 SEMESTER  Reference  Date Stamp 3 PRINCIPAL APPLICANT Surname:  Title:  Initials: Post held: Address: Telephone:  Fax: E-mail:  Is the applicant a possible observer? 4 COLLABORATORS Name:  Institute:  Observer? 5 SHORT TITLE OF PROPOSAL (maximum 12 words) 6 SUMMARY OF PROPOSED OBSERVATIONS 7 FOCAL STATION, INSTRUMENT AND DETECTOR Focal station:  Instrument:  Detector(s):  Gratings/Filters: AAT applicants must give full details on the appropriate instrument scheduling form 8 OBSERVING TIME REQUESTED THIS SEMESTER Time requested this semester  Dark:  Grey:  Bright:  specify nights Minimum useful allocation this semester  Dark:  Grey:  Bright:  or weeks: UKIRT applicants requiring dark time must justify this in section 16 9 COMPLETE THIS SECTION ONLY IF THIS IS A LONG TERM PROPOSAL Total time requested  Dark:  Grey:  Bright:  specify nights or weeks: Justification for long term status must be given in section 15 1.

INT

2001A

14 N

10

N

McMahon  Dr  R.G Lecturer/Royal Society Research Fellow Institute of Astronomy Madingley Road Cambridge CB3 OHA 01223-337548 01223-337523 rgm@ast.cam.ac.uk Yes P. Hewett, C. Mackay, I. Parry, C. Sabbey  Institute of Astronomy  Yes S. Hodgkin, M. Irwin, J. Lewis  Astronomical Survey Unit, Institute of Astronomy  Yes N. Walton Isaac Newton Group A Public Near IR imaging survey on the INT We propose to continue the major public public wide field near IR imaging survey we started in semester 2000A on the INT using the Cambridge Infra Red Survey Instrument(CIRSI). CIRSI is an innovative JH-band, wide field, IR camera built by Mackay, McMahon and Parry using a mosiac of 4 Rockwell 1024^{2} HgCdTe arrays. With an instantaneous field of view of 15.6’x15.6’ CIRSI is the largest IR camera in existence. Our immediate goal is to survey in J and H band ∼15deg^{2} to J=20.5; H=19.1(∼3.5magnitudes fainter than 2MASS) with the aim of extending the survey to 50deg^{2} or to fainter limits in later semesters. The survey will sample a ’local’ volume ∼10% that of 2MASS, albeit 20 times fainter. The range of scientific programs feasible with the data range from searches for cool white dwarfs to the identification of serendipitous XMM X-ray sources and is also urgently required for the design of UKIRT WFCAM and VISTA surveys.

Prime  CIRSI  4 Rockwell 1024^{2}  J,H

10 SCHEDULING INFORMATION  Preferred dates: Impossible dates: Give justification for impossible dates If observations are to be simultaneous with other telescopes or satellites, give details: Any other scheduling constraints: Include likely clashes with other time applications, constraints on lunar position or quarter, instrument preparation requirements, etc 11 SERVICE OBSERVING  yes:  no:  maybe: 12 SUPPORT ASTRONOMER REQUESTED AT TELESCOPE every night:  no:  first night only: 13 LIST OF PRINCIPAL TARGETS Object(s):  RA(h,m):  Dec(degs):  Mag(type):  Colour:  Exp. Time: 14 OTHER APPLICATIONS FOR OBSERVING TIME FOR THIS OR SIMILAR PROGRAMMES IN THIS SEMESTER Include all other telescope and satellite applications whether via PATT or any other time assignment committee Telescope/satellite:  Title of programme:  2.

May-June Feb-Mar Instrument unavailable

x

x

Fields  RA (J2000) Dec  Field size  bands Exp. Time notes
selected from  (other data)
NGC Equatorial Strip 10 < α <15 ±0.25  75^{∘}x0.5^{∘}  J,H  300secs  FIRST, NVSS, 2DF
SGC Equatorial Strip 22 < α <03 ±0.25  75^{∘}x0.5^{∘}  J,H  300secs  FIRST
WFSJ1610+54  16 10.0  +54 30  3^{∘}x3^{∘}  J,H  300secs  ISO-ELAIS, FIRST
WFSJ1637+41  16 37.0  +41 16  3^{∘}x3^{∘}  J,H  300secs  ISO-ELAIS, FIRST
XMM fields  all RAs  δ > −10 0.5^{∘}x0.5^{∘} each J,H  900secs  FIRST/NVSS

INT  XMM related WFS proposal by Watson
INT  WFS proposal by McMahon

15 SCIENTIFIC JUSTIFICATION Case not to exceed this A4 page, but one side of diagrams and/or references may be attached if desired 3.

Introduction Astronomy is an observational science, and quantitative surveys form a basic foundation of the subject. Without systematic, quantitative, unbiased surveys our subject will stagnate. The strategic value of such surveys is epitomised in plans for the UKIRT Wide Field upgrade and the VISTA JIF award. However, these programs are many years off. Recognising the obvious need for wide field IR surveys, the Institute of Astronomy, Cambridge has invested £600k in the development of CIRSI an innovative wide field JH-band IR camera built around a mosaic of 4 Rockwell 1024^{2} HgCdTe arrays. Currently most of the year CIRSI is based at the DuPont 2.5m in Chile, where the scientific emphasis is on deep IR surveys we are also keen to devote some time to the Northern Hemisphere and to carry out shallower surveys. Rather than embark on piecemeal small proposals, we propose to carry out a major moderate depth IR survey on the INT. CIRSI has now reached a stable operating level and it is therefore timely to embark on a major wide field public IR survey where reduced data products are made public via the ING archive on a 3-6month timescale after observations take place. This survey shall be carried out in regions that have extant optical CCD data from the INT WFS project(McMahon etal, 1999) and in regions surveyed by the XMM serendipitous survey(Watson etal; see WFS proposal). Below we summarise some science themes that are the focus of the proposers.

Cool White Dwarfs Microlensing results suggest that 50% of the Halo dark matter may be explained in 0.5 M_{⊙} objects (with a number density of 0.0079 pc^{3}, Richer et al. 1999). Cool white dwarfs are an obvious candidate, especially following the calculation of new model atmospheres (Hansen 1998) which show that they are blue in colours such as I–H, but red in g–r. We will search for objects with this trademark signature (Hodgkin et al. 1999), and provide a direct measurement of the space density of cool white dwarfs. If they are to explain the Halo dark matter then we expect to find ∼ 20 cool white dwarfs in our survey. Figure 1 shows the optical-IR spectrum of the cool white dwarf WD0346+46. Note the distinctive optical and IR colours.

Low mass stars and brown dwarfs Brown dwarfs (BDs) are the most numerous stars in the Galaxy, although their mass contribution is only around 10% (Reid et al. 1999). One of the most remarkable discoveries of the last year has been the almost simultaneous discoveries by DENIS, 2MASS and Sloan surveys of the first local BDs with the discovery of 5 T (Methane) dwarfs (T_{e} ≤ 1200 K, Burgasser et al. 1999) and 25 L dwarfs (1200 ≤ T_{e} ≤ 2000K, Reid et al. 1999). However the remarkable detection of a T dwarf in the ESO deep imaging survey within a volume of only 1 pc^{3} (Cuby et al. 1999) suggests a density 100 times higher for T dwarfs. With this proposal we can find T dwarfs out to >100 pc and test the ESO result. If the very local 2MASS volume is representative, then we will find ∼ 25 T dwarfs. If it transpires that the local volume is biased (as implied by the ESO result) then we will find many more. We shall also be able to search for L dwarfs later than L8, and T dwarfs cooler than Gliese 229B (T_{e} ∼ 900K; Figure 2).

Reddening independent selection of quasars and the neutral gas content of the Universe

Storrie-Lombardi, McMahon and Irwin(1996a) have recently shown via the study of damped Lyman-α absorption systems(log(N(HI)>20.3) superposed on the spectra of distant quasars that over the redshift range 2<z<4, the total mass in neutral gas is comparable with the total visible mass in stars in present day galaxies and that the volume density is constant over the redshift range 2 to 4( see Figure 3). The simplest interpretation is that over the redshift range 2 to 4, we are observing the assembly of disk galaxies. However, the use of quasar absorption lines in this manner has been severely critised by Fall & Pei(1995) since optical samples are biased against lines of sight that may contain gas rich dusty galaxies. The only way to unambiguously ascertain whether there is a problem in such work is to select quasars at IR wavelenghts. Figure 4 shows how we can select quasars using optical and IR colours analogous to the so called KX method for producing K-band flux-limited samples of quasars (Warren, Hewett & Foltz, 2000) The predicted surface density of QSOs with z>2 and H<18.5 is 5-10deg^{2} and thus in 50deg^{2} we expect 250-500 QSOs with z>2.

16 TECHNICAL INFORMATION (I) Give details of the technical feasibility of the proposal (S/N,etc) AND any non-standard technical requirements 4.

Table 1 - CIRSI Characteristics on 2.5m INT

Pixel scale (arcsec/pixel)  0.46
FOV per chip  7.8’ x 7.8’
effective FOV of mosaic  15.6’ x 15.6’
FOV for 2x2 mosaic  29.6’ x 29.6’
Frames/deg^{2}  16

Sensitivity( 5σ in 1.2” seeing)

J saturation time  90secs
H saturation time (sec)  45secs
J: 5mins  20.6
H: 5 mins  19.1

Table 2 - DENIS + 2MASS Characteristics

2MASS DENIS  INT  notes
Telescope diameter(m)  1.3  1.0  2.5m
Exposure time  7.8secs 10.0secs  300secs
Detectors  3 x 256^{2} 2 x 256^{2}  4 x 1024^{2}
Pixel Size(arcsecs)  2x2  3x3  0.46 x 0.46
Field of view  8.5’x8.5’ 12.8’x12.8’ 15.6’x15.6’

Magnitude limits(5sigma)

z  -  -  22.1  INT WFS
J  16.5  16.0  20.6
H  15.8  -  19.1
K  15.0  14.0  -

Survey Parameters The survey design requirement are primarily driven by the requirement that all objects are within reach of current generation optical or near IR low resolution spectrographs on 4-8m class telescopes. The survey area is tempered by the requirement that we can process and deliver processed data within 3-6months of the data being acquired and by the requirement that statistically significant samples can be discovered in each of the science areas. H band is primarily required for the quasar program since we want to minimize the effect of reddening. The J band data is primarily driven by the low mass star program. Time requirements Assuming 300secs per pointing split into 5 discrete dither positions and overheads for dithering, slewing, preread of 40%, it will take 500secs per pointing and 8000secs to cover 1deg^{2}. In a typical May/June night of 7.0hrs(time between twilight on 2001, June 1) we can survey 3.1deg^{2}. Thus in 7 nights, with no allowance for bad weather we can cover 22.0deg^{2} to a H=19.1(5sigma) or J=20.6 in seeing of 1.2” the current INT 75% quartile value. We are therefore requesting 10nights in 2001A to survey 15deg^{2} in both bands. In 2000B we shall request time to either inrease the area or to survey deeper in the same area. In addition we propose to spend 4 nights imaging 4 fields that have XMM data with exposure times 900secs per pointing. [see the WFS proposal from Watson etal] Survey Area We plan to carry out the survey centred predomintly on the WFS survey zero declination strip and 1610+41 field. ∼25% of the survey will be carried out in survey fields that will be covered by XMM. The proposed allocation is constained by the availability of CIRSI on the INT and the amount of data that we believe we can process in a timely manner taking into account the requirement to also process the INT WFC data during the same period. All the survey areas are covered by the FIRST radio survey. Long Term Strategy This proposal is only asking for time in semester 2001A. Normally CIRSI spends the B semester in Chile but we are considering bringing it to the ING for semester 2000B so that progress on this survey program can be accelerated. A decision on this will be made in Nov, 2000. We plan to put in a more substantial proposal at the next deadline where we shall ask for time to complete the survey.

17 SUMMARY OF BACKUP PROGRAMME FOR POOR OBSERVING CONDITIONS If instrumentation or setup differs from main programme, give full details 18 RELATED PATT APPLICATIONS OVER THE LAST FOUR SEMESTERS (including unsuccesful applications) PATT reference:  Award:  Clear nights: Comments: 19 PUBLICATIONS BASED ON PATT TIME PUBLISHED DURING THE LAST FOUR SEMESTERS (maximum 6) 20 EXPERIENCE OF INTENDED OBSERVERS WHO HAVE NOT PREVIOUSLY USED THIS TELESCOPE 21 COMPLETE IF THE OBSERVATIONS ARE PRIMARILY FOR A STUDENT RESEARCH TRAINING PROGRAMME Name of student: Project title: 22 COMPLETE IF THE OBSERVATIONS ARE ASSOCIATED WITH A CURRENT PPARC/SERC RESEARCH GRANT Name of principal investigator: Grant title: Grant number: 23 NON-STANDARD TRAVEL AND SUBSISTENCE REQUIREMENTS (UK observers only) Justify requests for travel and subsistence for more than one person: Details of any other expenditure (eg freight, remote observing): 5.

Observations can be carried out in non-photometric conditions via the use of overlapping exposures

W/99A/56 3Bn 80% useable Paper in press; Gray etal(astro-ph/astro-ph/0004161; see below)

I/00A/26 7Bn 40% useable Observations were in June 2000; data being reduced; 5-7deg^{2} surveyed in H band; 1-2 deg^{2} in J(data quality uncertain) This is a resubmission

(i) ’High redshift quasars in the INT Wide Field Survey’, Sharp, McMahon, Irwin & Hodgkin 2000 in prep; (ii) ’Infrared observations of gravitational lensing in Abell 2219 with CIRSI’ with Gray et al, 2000, MNRAS in press(astroph/astro-ph/0004161) (iii) ’The PSCz catalogue’, with Saunders etal, 2000, MNRAS, 317. (iv) ’Measurements of Omega and Lambda from 42 High-Redshift Supernovae with Perlmutter,S. et al, 1999, ApJ, 517, 565;

(v) ’Submillimetre observations of luminous z>4 radio-quiet quasars and the contribution of AGN to the submm source population’, McMahon, Priddey, R, et al, 1999, MNRAS, 309, pp1. (vi) ’The power spectrum of the Point Source Catalogue redshift survey’, with Sutherland et al, 1999, MNRAS, 308, 289

Dr P C Hewett

PATT related travel and subsistence 2000-2004 PPA/G/O/1999/00581

CIRSI is a private non-common-user instrument and installation, astronomical operation and technical support is provided by the IOA Instrument team. We therefore we request PATT funding for a second observer as specified in our PATT rolling grant.

Shipping to and from La Palma once per semester: £1800 as specified in our PATT rolling grant.

PATT PATT2 substyle file L^{A}TEX macros v0.1 03 Sep 96.